# Latent Action Diffusion for Cross-Embodiment Manipulation

### v2

@@ 4 Experimental Results and Discussion
We conducted experiments covering three different end-effectors and three tasks across two setups: one with the Faive hand and the Franka gripper and one setup with the mimic hand and the Franka gripper. For each end-effector in each setup, we compare single-embodiment policies with one cross-embodiment policy co-trained on data from all end-effectors. Additionally, we validate our design choices for our contrastive action model through an ablation study.
4.1 Ablation Study: Contrastive Action Model
@@
To validate our design choices, we compare several versions of the contrastive action model (Table 1). As metrics, we utilize self-reconstruction (SR) and cross-reconstruction (CR) validation losses. The ablation without temperature annealing keeps the temperature constant at the previous final value. The ablation without finetuning freezes the encoders in the second training step while the decoders are being trained. Both temperature annealing and finetuning the encoders reveal themselves to substantially improve both self- and cross-reconstruction metrics, with finetuning being the most important addition to the pipeline.
(a) Pick-and-place of a plush toy into a bowl in different locations with the Faive hand and Franka gripper.
(b) Pick-and-place of a plastic cube into a box in different locations with the mimic hand and Franka gripper.
Figure 5: We compare policies co-trained on cross-embodiment data (ours) with single-embodiment policies trained using both latent and explicit actions. For all policies, we show the mean and standard deviation of successful task executions across all object locations. The task setups are shown in Figure 7.
4.2 Setup #1: Pick and Place with Faive Hand and Franka Gripper
We train policies with data collected using a Franka parallel gripper and a dexterous Faive hand for picking and placing a plush toy inside a bowl. For each end-effector, we collected 100 episodes of data, varying the position of the target object and the place location. The latent action space for this setup was learned as shown in Figure 2 using human hand data, Faive hand data, and data from the Franka gripper. For both end-effectors, we only utilize a single external RGB camera as observation. For evaluation, we collect 40 grasps for each embodiment, consisting of 5 tries across 8 grasp locations.
4.2.1 Performance
The performance for the different policies is shown in Figure 5(a). Single-embodiment policies with and without latent actions show no noticeable performance difference. The cross-embodiment policy outperforms both single-embodiment policies by 10% and 7.5% respectively. The reduced standard deviation shows that the grasp success is more consistent across different grasping positions.
4.3 Setup #2: Pick and Place with mimic Hand and Franka Gripper
We train policies with data collected using the mimic hand and the Franka gripper for placing a plastic cube inside a box. For each embodiment, we collect 250 episodes of data. The latent space was learned with data from the mimic hand and the Franka gripper. In addition to an external camera view, we utilize the pose of the Franka arm as observations as well as a wrist camera view for the mimic hand, which is replaced by zero-padding for the Franka gripper. For evaluation, we collect 18 grasps for each embodiment, consisting of 3 tries across 6 grasp locations.
4.3.1 Performance
The performance for the different policies is shown in Figure 5(b). Single-embodiment policies with and without latent actions show a slight increase in performance for the mimic hand and no increase in performance for the Franka gripper. The cross-embodiment policy for the Franka gripper underperforms the respective single-embodiment policies with a reduction in performance (we attribute this to the lacking wrist camera and incidental factors during data collectionz). In contrast, the cross-embodiment policy rolled out for the mimic hand outperforms its single-embodiment equivalents with a 13% increase in success rates over the best single-embodiment policy.
4.4 Setup #3: Block Stacking with mimic Hand and Franka Gripper
We train policies with data from the mimic hand and the Franka gripper to grasp a block, stack it on top of another block by aligning the pins and then pick up and place the stacked blocks into the box (Figure 6(b)). We collect 200 demontrastions for each embodiment. Both policies use a single external camera and the pose of the Franka arm as observations. For evaluation, we perform 70 rollouts per policy per embodiment.
4.4.1 Performance
We show the completion rates for the individual task stages in Figure 6. For the first stage, a coarse manipulation skill, our method exhibits strong performance with absolute completion rate improvements of 25% and 13%. For the fine-grained manipulation stage, performance slighly decreases for the mimic hand, likely due to the hand partially occluding the object from the camera. For the Franka gripper, the performance improves significantly by 13% and 11%.
4.5 Discussion
Latent actions accurately encode diverse action spaces. When comparing single-embodiment latent policies to their counterparts with explicit actions, there is no significant performance decrease, which suggests that the learned latent spaces can accurately encode the action spaces of the three embodiments with which we conduct the experiments.
Latent action representations enable multi-robot control and cross-embodiment skill transfer. With our methodology, a single policy can control two highly different end-effectors. Furthermore, latent policies learn helpful shared representations that we observe in improved success rates when co-training on data from different end-effectors for both coarse and fine-grained manipulation tasks. We attribute the performance drop of the Franka gripper in setup #2 to the asymmetric observations: co-training with missing camera views remains an open challenge (similar to [3]). Further, the slight performance decrease for the mimic hand in setup #3 is likely due to the hand occluding the block after grasping it, making fine-grained control more difficult.
In summary, our approach successfully demonstrates that latent action spaces can unify control across diverse robotic embodiments while enabling improved performance through cross-embodiment skill transfer. Learning with asymmetric observation spaces remains remains an important challenge for scaling policy learning.
(a) Task stage completion rates for each task stage for single-embodiment diffusion policy versus cross-embodied latent diffusion policy (ours).
(b) Task stages: i) initial setup ii) picking the first block iii) stacking/inserting it on top of the second block iv) putting both into the box.
Figure 6: We evaluate our methodology across 70 trials per policy per embodiment on a highly challenging block stacking task, comprised of three stages. Both the mimic hand and Franka gripper exhibit a significant performance gain for coarse-grained manipulation (stage 1), whereas performance for fine-grained manipulation is largely improved for the Franka gripper.
5 Conclusion
Among current challenges in robotics, enabling effective skill-transfer across diverse embodiments is of crucial importance to both maximize the volume and diversity of suitable training data and to ensure the reusability of training data throughout the lifecycle of different end-effectors. To this end, we frame cross-embodiment learning with different end-effectors as a multimodal representation learning problem and propose a four-stage pipeline to learn capable policies that can control multiple end-effectors. In real-world experiments with the dexterous hands and a Franka parallel gripper, we demonstrate that through co-training on cross-embodiment data with our method for latent action spaces, both multi-robot control and positive skill transfer across embodiments are possible. In particular, the performance improvement of up to 25% (average: 13.4%) indicates that our method facilitates skill transfer between end-effectors with a large embodiment gap and underlines its potential for wider use across a broader range of robot morphologies. Future work includes expanding our method to a more diverse ecosystem of end-effectors and further investigating the behavior of skill transfer across different dataset sizes with more distinct visual differences.
6 Limitations
The most common failure mode is that the policy fails to position the wrist in a suitable position to grasp the object. The positioning of the end-effector is especially important for the comparatively small Franka gripper, which explains the consistently lower manipulation performance compared to the humanoid hands. If the wrist is correctly positioned, grasps are typically successful.
Asymmetric Dataset Sizes
We find that adding datasets such as BridgeV2 [23] or DexYCB [24] does not yet improve the performance of the policy. While their diverse respective action spaces can be unified through our method, visual differences and the highly asymmetric scale in dataset sizes still present significant challenges for achieving skill transfer via co-training.
@@ 6 Limitations
Skill transfer in the presence of asymmetric observations (e.g. one embodiment has an additional camera view) remains an open challenge for future work, which is reflected in our experiments.
Ambiguity in Action Space Mapping
The current contrastive learning method for learning joint action spaces does not automatically guarantee high quality reconstruction for all embodiments. Presently, it is crucial to empirically evaluate the encoders and decoders to verify that reconstruction and cross-reconstruction errors are low, before beginning policy training.
@@
Given that our method for contrastive latent space learning has no implicit latent space regularization (unlike VAE-based methods), there is a risk that the latent space might be suboptimally non-smooth and hard to model for the downstream policy. Future work is needed to guarantee the smoothness of the latent space while maintaining alignment and high reconstruction accuracy.
Future work: Larger scale experiments and more embodiments
Present results are still limited to small-scale experiments on a relatively low number of embodiments (albeit a higher diversity of end-effector morphology compared to other cross-embodiment works). We leave larger-scale experiments to future work.
Acknowledgments
We express our gratitude to Emanuele Palumbo for valuable discussions and insights on multimodal learning architectures. Furthermore, we are grateful for support from Chenyu Yang in setting up the initial version of srl_il and Davide Liconti in writing the control interface for the Franka gripper. Additionally, we thank the mimic team for support: Victor Montesinos and Jonas Pai for software, Benedek Forrai and Robert Malate for robotics, Philipp Wand for hardware, Stephan Polinski for electronics, and Norica Bacuieti for data collection.
We are grateful for grant funding from Innosuisse (122.679 SIP) and ESA BIC received by mimic robotics in connection to this project. We are also grateful to funding supporting academic co-authors: funding from the ETH AI Center, SNSF Project Grant MINT #200021_215489, Swiss Data Science Center, Amazon Research Award, Armasuisse.

### v3

@@ IV Experimental Results and Discussion
We conducted experiments covering three different end-effectors and three tasks across two setups: one with the Faive hand and the Franka gripper and one setup with the mimic hand and the Franka gripper. For each end-effector in each setup, we compare single-embodiment policies with one cross-embodiment policy co-trained on data from all end-effectors. We train policies using contrastive action models that were trained for the end-effectors present in each task. In addition, we validate our design choices for our contrastive action model through an ablation study.
IV-A Ablation Study: Contrastive Action Model
To validate our design choices, we compare several versions of the contrastive action model (Table III). As metrics, we utilize self-reconstruction (SR) and cross-reconstruction (CR) validation losses. The losses shown in the table are denormalized losses for the 16-dimensional action space of the mimic hand and the 1-dimensional action space of the Franka gripper. Due to the higher dimension and complexity of the larger action space, the losses for the mimic hand are larger than the losses for the Franka gripper.
@@ IV-A Ablation Study: Contrastive Action Model
We compare the full training pipeline for the contrastive model to three ablations. The ablation without temperature annealing keeps the temperature constant at the previous final value. The ablation without finetuning freezes the encoders in the second training step while the decoders are being trained. Both temperature annealing and finetuning the encoders reveal themselves to substantially improve both self- and cross-reconstruction metrics, with finetuning being the most important addition to the pipeline.
IV-B Experimental Setups
We evaluate our framework across three manipulation tasks using different combinations of end-effectors. For the Block Stacking task, we use the mimic hand and Franka gripper to pick a block, stack it on another, and place the pair into a box. For this task, we collected 200 demonstrations per embodiment using a single external camera for visual observations. Each policy is evaluated over 70 trials.
For the Block Pick & Place task, we use the mimic hand and the Franka gripper to place a plastic cube inside a box. We collected 200 demonstrations for each embodiment. In addition to an external camera and the robot arm pose, this setup uses a wrist camera for the mimic hand, which is replaced by zero-padding for the Franka gripper to test learning with asymmetric observations. Each policy is evaluated over 25 trials.
Finally, for the Plush Toy Pick & Place task, we use a Franka parallel gripper and a dexterous Faive hand to pick up a plush toy and place it into a bowl. For this setup, we collected 100 demonstrations for each end-effector and used a single external RGB camera for observations. Each policy is evaluated over 40 trials.
IV-C Results and Discussion
Our experiments demonstrate that co-training policies in a learned latent action space enable both multi-robot control from a single policy and significant performance gains through cross-embodiment skill transfer. We analyze the performance of our co-trained policies against single-embodiment baselines across three distinct manipulation tasks.
The multi-stage Block Stacking task showed the most pronounced skill transfer. For the initial coarse manipulation stage (”Pick block”), the co-trained policy yielded absolute success rate improvements of 25.3% for the mimic hand and 13% for the Franka gripper. More notably, the Franka gripper performance in the subsequent fine-grained stages (”Stack blocks” and ”Put into box”) improved significantly by 13% and 11%, respectively. This indicates that the gripper effectively learned more precise manipulation strategies from the dexterous hand data. The mimic hand saw a slight performance decrease in the final stage, likely due to the hand occluding the block from the camera after grasping it, making the final placement more difficult.
For the Block Pick & Place task, which featured asymmetric camera observations, the benefits of co-training were still evident for the more sensor-rich embodiment. The success rate of the mimic hand improved by 13% over its single-embodiment baseline. However, the Franka gripper’s performance decreased. We attribute this to the policy becoming reliant on the wrist camera view available only to the mimic hand, highlighting that skill transfer in the presence of asymmetric observations remains a significant challenge, despite the aligned action space.
In the Plush Toy Pick & Place task, the co-trained policy substantially outperformed the single-embodiment versions, achieving a 10% higher success rate for the Faive hand and a 7.5% improvement for the Franka gripper. This suggests that the policy learns a more robust, shared representation of the task, allowing skills learned from the dexterous hand (e.g., object handling) to benefit the simpler gripper, and vice versa.
In summary, our approach successfully unifies control across diverse robotic hardware and facilitates cross-embodiment skill transfer. The results confirm that policies learn helpful shared representations, leading to improved success rates in both coarse and fine-grained manipulation.
IV-D Limitations
@@
Our policy struggles to achieve consistent performance on all embodiments if they have different sensor modalities (e.g., wrist camera for one robot only). This remains a fundamental challenge for cross-embodiment learning. Future work could explore vision encoder adaptation or learned observation alignment.
@@
Our latent space is not explicitly regularized, which may lead to non-smooth regions that are harder to model for downstream policies. Integrating VAE-style priors or enforcing smoothness via time-contrastive losses could guarantee desirable latent space properties.
Dataset Scale Imbalance
We observe limited gains when adding large-scale datasets (e.g., BridgeV2 [23], DexYCB [24]) to our co-training mix. Training on datasets with highly imbalanced sizes and transferring skills despite significant environmental differences remains a fundamental challenge of cross-embodiment learning that could be explored via more sophisticated sampling strategies to determine dataset weights and visual representation learning that facilitates skill transfer given largely disjunct observation spaces.
V Conclusion
Among current challenges in robotics, enabling effective skill transfer across diverse embodiments is of crucial importance to both maximize the volume and diversity of suitable training data and to ensure the reusability of training data throughout the life cycle of different end-effectors. To this end, we frame cross-embodiment learning with different end-effectors as a multimodal representation learning problem and propose a two-stage pipeline to learn capable policies that can control multiple end-effectors. In real-world experiments with dexterous hands and a Franka parallel gripper, we demonstrate that through co-training on cross-embodiment data with our method for latent action spaces, we enable both multi-robot control and positive skill transfer across embodiments. In particular, the performance improvement of up to 25.3% (average: 13.4%) indicates that our method facilitates skill transfer between end-effectors with a large embodiment gap and underlines its potential for wider use across a broader range of robot morphologies. Future work includes expanding our method to a more diverse ecosystem of end-effectors and further investigating the behavior of skill transfer across different dataset sizes with more distinct visual differences.